# Unsupervised Learning of Effective Quantum Impurity Models

### v2

@@ Abstract
In particular, a prerequisite for the design and use of complex quantum nanoelectronics devices with advanced functionality beyond the classical paradigm, is a fundamental understanding of their low-temperature correlated electron physics and quantum transport properties. However, this is a notoriously difficult theoretical challenge because of the subtle interplay between the orbital and spin complexity of the nanostructure, determined by its structure and chemistry; strong electron interactions due to quantum confinement; and the coupling to $\sim 10^{23}$ conduction electrons in the external circuit. This results in nontrivial quantum phenomena such as Coulomb blockade [40], various forms of Kondo effect [40, 41, 38], and quantum interference [42, 43] – all of which strongly affect low-temperature electronic conductance through the device, and hence its functionality. As with coupled quantum dot devices [44, 45, *mitchell2016universality, 47], entangled spin and charge degrees of freedom can give rise to new physics in single molecule junctions. It is therefore a formidable task to derive simplified effective models that can still describe this range of physics in these kinds of quantum device.
@@
In the context of molecular electronics, the complex microscopic models describing single-molecule junctions can be mapped to much simpler effective two-channel Kondo models using a perturbative approach, as described in Refs. [38, 43]. This method captures simultaneously the effect of quantum interference and Kondo physics – but is inevitably approximate due to the perturbative treatment. Since thermodynamic observables flow under RG and the bare and effective model are defined at different energy scales, it is also not a priori clear whether such minimal Kondo models are sufficiently general to reproduce local observables of interest in the bare model [48]. In this paper we show how an RG analysis of effective interaction terms can be used to determine thermodynamic observables that are comparable across different energy scales. This allows us to introduce a novel machine learning (ML) methodology to derive accurate effective models for complex quantum impurity problems, that works by optimizing generalized (minimally constrained) models, and ensuring that local observables are correctly reproduced. We show by way of explicit examples that the low-energy Kondo physics is simultaneously captured. The parameters of the effective model are optimized by minimizing the Kullback-Leibler divergence (KLD) [49] that compares its ensemble of thermal states to that of the bare model. Information on the target is extracted from a numerical simulation of the system of interest. However, the full solution of the bare model is not required: minimization of the KLD requires only an estimation of thermal expectation values corresponding to specific local operators at relatively high temperatures. This can be achieved with any suitable quantum impurity solver [50, 51, 52, 53, *stadler2016interleaved]. Furthermore, the KLD can be shown to be convex under reasonable assumptions (see Appendix A) and its gradient is known analytically in closed form [55], which makes it an ideal optimization problem. We refer to this method as ‘unsupervised model learning’ (UML). We demonstrate the efficacy of the UML method by application to the Anderson impurity model (AIM) [36] and obtain some new non-perturbative results for this old problem. Finally we give an outlook to the application of this framework to more complex problems, where the development of tractable effective models is essential for the study of nontrivial correlated electron physics at low temperatures.
@@ I Overview of the Method
$D_{KL}\left[\hat{\rho}_{2}:\hat{\rho}_{1}\right]=\operatorname{tr}\left[\hat{\rho}_{1}\log(\hat{\rho}_{1})\right]-\operatorname{tr}\left[\hat{\rho}_{1}\log(\hat{\rho}_{2})\right]\;.$ (2)
The generalized KLD quantifies how distinguishable $\hat{\rho}_{1}$ is from $\hat{\rho}_{2}$. The thermal density matrix $\hat{\rho}=\frac{1}{\mathcal{Z}}e^{-\beta\hat{H}}$ is fully defined by the system Hamiltonian $\hat{H}$ (and inverse temperature $\beta$). Here $\mathcal{Z}={\rm tr}[e^{-\beta\hat{H}}]$ is the partition function of $\hat{H}$. Thus, the thermal density matrix can be seen as a proxy for its defining Hamiltonian, and the KLD as a measure of distinguishability between two Hamiltonians. To emphasize this we denote the KLD for two thermal density matrices $\hat{\rho}_{1}=\frac{1}{\mathcal{Z}_{1}}e^{-\beta\hat{H}_{1}}$ and $\hat{\rho_{2}}=\frac{1}{\mathcal{Z}_{2}}e^{-\beta\hat{H}_{2}}$ as $D_{KL}[\hat{H}_{2}:\hat{H}_{1}]$.
@@
The optimization problem is efficient, given the convexity of Eq. 9. The computational cost of the optimization is therefore mainly controlled by the method used to compute the observables comprising the gradient in Eq. 10. The bottleneck is then the calculation of the observables $\langle\hat{h}_{i}\rangle_{\rm bare}$ in the bare model. However, this needs to be done just once (e.g. with quantum Monte Carlo (QMC) methods [59, 50]). On the other hand, although multiple calculations of $\langle\hat{h}_{i}\rangle_{\rm eff}$ are required during the optimization, these are performed on the simplified effective model and are therefore by construction inexpensive (they can be done with e.g. the numerical renormalization group (NRG) method [51, 52, 53, *stadler2016interleaved]). For the results presented in the following, we have used continuous-time QMC (CT-QMC) and NRG to calculate the observables and the Adagrad or Adam GD methods [60].
@@ III.2.2 Optimizing the set of matched observables
As noted above, there is some ambiguity in Eq. 42 as to which projector to eliminate in the UML observable matching process. In Eq. III.1 we somewhat arbitrarily eliminated the projector $\hat{P}_{0,0}$ by using the identity operator, but in principle $\hat{P}_{0,1}$ or $\hat{P}_{\pm 1,1/2}$ could have been eliminated. Here we present a method for identifying which projector should be eliminated, and validate this numerically the comparing effective models based on different choices. For the case of the MC effective model it turns out that indeed Eq. III.1 is the optimal model. However for the more complicated effective models considered in Sec. IV which feature a larger number of observables, one requires a systematic approach
@@
The first step is to define the Frobenius scalar product (FSP) on the operator space $\langle\hat{A},\hat{B}\rangle={\rm tr}\left[\hat{A}^{\dagger}\cdot\hat{B}\right]$ and the corresponding norm $\|\hat{A}\|=\sqrt{\langle\hat{A}^{\dagger},\hat{A}\rangle}$. Using the FSP any operator $\hat{h}_{i}$ can be decomposed into its contributions from the low-energy fixed point Hamiltonian of the bare model $\hat{H}^{*}$, and RG-irrelevant corrections to the fixed point, $\hat{\Lambda}_{i}$. We define $\alpha_{i}=\langle\hat{h}_{i},\hat{H}^{*}\rangle/\|\hat{H}^{*}\|^{2}$ and $\beta_{i}=\langle\hat{h}_{i},\hat{\Lambda}_{i}\rangle/\|\hat{\Lambda}_{i}\|^{2}$, noting that $\langle\hat{\Lambda}_{i},\hat{H}^{*}\rangle=0$. The operator $\hat{\Lambda}_{i}$ can be found by the Gram-Schmidt procedure $\hat{\Lambda}_{i}=\hat{h}_{i}-\alpha_{i}\hat{H}^{*}$. We now decompose the projector as $\theta_{i}\hat{h}_{i}=\theta_{i}(\alpha_{i}\hat{H}^{*}+\beta_{i}\hat{\Lambda})\equiv\gamma_{i}\hat{H}^{*}+\lambda_{i}\hat{\Lambda}_{i}$, which can be used for the substitution,
@@ IV.3 UML optimization of extended models
Figure 6: Analysis of UML optimization of the extended MC effective models (Eqs. 59 and 61), taking the AIM as our bare model. (a) Comparison of relative errors in computed observables (Eq. 63) as a function of the GD optimization parameter update step (epoch). Lines are color-coded according to the key, labelled by the (diagonal) projector quantum numbers $(Q,S,T)$. (b) Convergence of the Kondo temperature $T_{K}$ with epoch (red line) compared with the reference result in the bare model (blue line). (c) Comparison of relative (logarithmic) error in the Kondo temperature for the fixed-basis effective model Eq. 59 (purple line) and generalized basis model Eq. 61 (red line). Plotted for $0.38V^{2}/U_{d}=$ 0.3 and $U_{d}=0.4$, with the optimization performed at $T=0$ using NRG as the impurity solver for both bare and effective models.
@@
Taking the AIM again as our bare model for demonstration, we perform UML optimization on the extended MC effective model Eq. 61. For representative AIM parameters, in Fig. 6(a) we plot the relative error of selected (diagonal) observables labelled by the projector quantum numbers $(Q,S,T)$, as a function of the optimization step (epoch). We find that most effective model observables converge exponentially quickly to their reference bare model counterparts, whereas a few are essentially stationary and cannot be matched. However, as Fig. 6(b) shows, this does not make much difference to the quality of the optimized effective model, as quantified by the Kondo temperature $T_{K}$. After $\sim 100$ or so training steps, $T_{K}$ for the effective model (red line) converges quite accurately to that of the bare model (blue line).

### v3

@@ Abstract
A prerequisite for the design and use of complex quantum nanoelectronics devices with advanced functionality beyond the classical paradigm (for example in quantum metrology [40, 41, *mihailescu2023thermometry]), is a fundamental understanding of their low-temperature correlated electron physics and quantum transport properties. However, this is a notoriously difficult theoretical challenge because of the subtle interplay between the orbital and spin complexity of the nanostructure, determined by its structure and chemistry; strong electron interactions due to quantum confinement; and the coupling to $\sim 10^{23}$ conduction electrons in the external circuit. This results in nontrivial quantum phenomena such as Coulomb blockade [43], various forms of Kondo effect [43, 44, 38], and quantum interference [45, 46] – all of which strongly affect low-temperature electronic conductance through the device, and hence its functionality. As with coupled quantum dot devices [47, 48, *mitchell2016universality, 50], entangled spin and charge degrees of freedom can give rise to new physics in single molecule junctions. It is therefore a formidable task to derive simplified effective models that can still describe this range of physics in these kinds of quantum device.
@@
In the context of molecular electronics, the complex microscopic models describing single-molecule junctions can be mapped to much simpler effective two-channel Kondo models using a perturbative approach, as described in Refs. [38, 46]. This method captures simultaneously the effect of quantum interference and Kondo physics – but is inevitably approximate due to the perturbative treatment. Since thermodynamic observables flow under RG and the bare and effective model are defined at different energy scales, it is also not a priori clear whether such minimal Kondo models are sufficiently general to reproduce local observables of interest in the bare model [51]. In this paper we show how an RG analysis of effective interaction terms can be used to determine thermodynamic observables that are comparable across different energy scales. This allows us to introduce a novel machine learning (ML) methodology to derive accurate effective models for complex quantum impurity problems, that works by optimizing generalized (minimally constrained) models, and ensuring that local observables are correctly reproduced. We show by way of explicit examples that the low-energy Kondo physics is simultaneously captured. The parameters of the effective model are optimized by minimizing the Kullback-Leibler divergence (KLD) [52] that compares its ensemble of thermal states to that of the bare model. Information on the target is extracted from a numerical simulation of the system of interest. However, the full solution of the bare model is not required: minimization of the KLD requires only an estimation of thermal expectation values corresponding to specific local operators at relatively high temperatures. This can be achieved with any suitable quantum impurity solver [53, 54, 55, 56, *stadler2016interleaved]. Furthermore, the KLD can be shown to be convex under reasonable assumptions (see Appendix A) and its gradient is known analytically in closed form [58], which makes it an ideal optimization problem. We refer to this method as ‘unsupervised model learning’ (UML). We demonstrate the efficacy of the UML method by application to the Anderson impurity model (AIM) [36] and obtain some new non-perturbative results for this old problem. Finally we give an outlook to the application of this framework to more complex problems, where the development of tractable effective models is essential for the study of nontrivial correlated electron physics at low temperatures.
@@ I Overview of the Method
$D_{KL}\left[\hat{\rho}_{1}:\hat{\rho}_{2}\right]=\operatorname{tr}\left[\hat{\rho}_{1}\log(\hat{\rho}_{1})\right]-\operatorname{tr}\left[\hat{\rho}_{1}\log(\hat{\rho}_{2})\right]\;.$ (2)
The generalized KLD quantifies how distinguishable $\hat{\rho}_{1}$ is from $\hat{\rho}_{2}$. The thermal density matrix $\hat{\rho}=\frac{1}{\mathcal{Z}}e^{-\beta\hat{H}}$ is fully defined by the system Hamiltonian $\hat{H}$ (and inverse temperature $\beta$). Here $\mathcal{Z}={\rm tr}[e^{-\beta\hat{H}}]$ is the partition function of $\hat{H}$. Thus, the thermal density matrix can be seen as a proxy for its defining Hamiltonian, and the KLD as a measure of distinguishability between two Hamiltonians. To emphasize this we denote the KLD for two thermal density matrices $\hat{\rho}_{1}=\frac{1}{\mathcal{Z}_{1}}e^{-\beta\hat{H}_{1}}$ and $\hat{\rho_{2}}=\frac{1}{\mathcal{Z}_{2}}e^{-\beta\hat{H}_{2}}$ as $D_{KL}[\hat{H}_{1}:\hat{H}_{2}]$.
@@
The optimization problem is efficient, given the convexity of Eq. 9. The computational cost of the optimization is therefore mainly controlled by the method used to compute the observables comprising the gradient in Eq. 10. The bottleneck is then the calculation of the observables $\langle\hat{h}_{i}\rangle_{\rm bare}$ in the bare model. However, this needs to be done just once (e.g. with quantum Monte Carlo (QMC) methods [62, 53]). On the other hand, although multiple calculations of $\langle\hat{h}_{i}\rangle_{\rm eff}$ are required during the optimization, these are performed on the simplified effective model and are therefore by construction inexpensive (they can be done with e.g. the numerical renormalization group (NRG) method [54, 55, 56, *stadler2016interleaved]). For the results presented in the following, we have used continuous-time QMC (CT-QMC) and NRG to calculate the observables and the Adagrad or Adam GD methods [63].
@@ 0 $1$ $0,~{}\pm 1$
In terms of the interpretability of effective models learned by UML, we note that the symmetry analysis already offers insights into the kinds of coupling terms that can arise in a given extended impurity space. The fact that the physical impurity hosts a local moment is exploited directly to reduce the dimensionality of the active impurity space; this has implications for the interactions appearing in the effective model. By formulating the resulting symmetry projectors in second-quantized form, as per Eq. III.1, we obtain a more physical interpretation of these terms. Indeed, recombining these in the form of Eq. 47 makes a connection to the RG theory of the low-energy properties of the bare model. However, the numerical values of the optimized coupling constants remain somewhat inscrutable. Only in perturbative limits do these effective parameters take a simple form in terms of the bare model parameters.
@@ III.2.2 Optimizing the set of matched observables
As noted above, there is some ambiguity in Eq. 42 as to which projector to eliminate in the UML observable matching process. In Eq. III.1 we somewhat arbitrarily eliminated the projector $\hat{P}_{0,0}$ by using the identity operator, but in principle $\hat{P}_{0,1}$ or $\hat{P}_{\pm 1,1/2}$ could have been eliminated. Here we present a method for identifying which projector should be eliminated, and validate this numerically by comparing effective models based on different choices. For the case of the MC effective model it turns out that indeed Eq. III.1 is the optimal model. However for the more complicated effective models considered in Sec. IV which feature a larger number of observables, one requires a systematic approach
@@
The first step is to define the Frobenius scalar product (FSP) on the operator space $\langle\hat{A},\hat{B}\rangle={\rm tr}\left[\hat{A}^{\dagger}\cdot\hat{B}\right]$ and the corresponding norm $\|\hat{A}\|=\sqrt{\langle\hat{A}^{\dagger},\hat{A}\rangle}$. Here we use the FSP to decompose an operator $\hat{h}_{i}$ from the effective model into contributions from the low-energy fixed point Hamiltonian of the bare model $\hat{H}^{*}$, and contributions from RG-irrelevant corrections to the fixed point, $\hat{\Lambda}_{i}$. We define $\alpha_{i}=\langle\hat{h}_{i},\hat{H}^{*}\rangle/\|\hat{H}^{*}\|^{2}$ and $\beta_{i}=\langle\hat{h}_{i},\hat{\Lambda}_{i}\rangle/\|\hat{\Lambda}_{i}\|^{2}$, noting that $\langle\hat{\Lambda}_{i},\hat{H}^{*}\rangle=0$. The operator $\hat{\Lambda}_{i}$ can be found by the Gram-Schmidt procedure $\hat{\Lambda}_{i}=\hat{h}_{i}-\alpha_{i}\hat{H}^{*}$. We now decompose the projector as $\theta_{i}\hat{h}_{i}=\theta_{i}(\alpha_{i}\hat{H}^{*}+\beta_{i}\hat{\Lambda})\equiv\gamma_{i}\hat{H}^{*}+\lambda_{i}\hat{\Lambda}_{i}$, which can be used for the substitution,
@@ III.2.4 UML for complex impurities
Ultimately, we envision that UML will be integrated with ab initio methods to tackle problems that are currently out of reach for brute-force calculations. This is beyond the scope of the present work, but we provide a discussion of how few-body approximants to bare-model observables might play a role in UML in Appendix D.
Figure 6: (a) Schematic of the triple quantum dot bare model (blue) and the associated MC effective model (orange). (b) Entropy $S_{\rm imp}(T)$ vs $T/B$ for the TQD bare model (dot-dashed lines) and the UML-optimized effective model (solid lines), for different magnetic field strengths $B$. (c) Corresponding behavior of the magnetic susceptibility $T\chi_{\rm imp}(T)$. Plotted for TQD model parameters $U_{d}=1.1$, $V=0.25$, $t=0.10$, $t^{\prime}=0.15$, $\zeta=0.05$, $D=1$. Both bare and effective models solved using NRG. Quantitative agreement is found in all cases in the regime $T\ll U_{d}$ where the mapping is expected to hold.
III.2.5 Application: triple quantum dot device
As a nontrivial demonstration of the UML methodology, we now consider its application to a bare model describing a triangular triple quantum dot (TQD) coupled to a metallic lead, in a magnetic field – see Fig. 6(a). This system supports a rich range of physics [81, *mitchell2013local, 83, 84]. The physical impurity Hilbert space has dimension $d_{\mathcal{H}}=4^{3}=64$, and SU(2)-spin symmetry as well as particle-hole symmetry is broken. We consider a regime of filling in which the TQD hosts a net spin-$\tfrac{1}{2}$ moment, delocalized across the three interacting quantum dot sites. The bare model Hamiltonian is given by,
$\begin{split}&\hat{H}^{\text{TQD}}_{\text{bare}}=\hat{H}^{\text{bath}}_{\rm disc}+\tfrac{1}{2}U_{d}\sum^{3}_{i=1}(\hat{n}^{d}_{i}-1)^{2}+\zeta\hat{N}^{d}+B\hat{S}^{d}_{z}\\ &+\sum_{\sigma}\big{[}t\hat{d}^{\dagger}_{1\sigma}(\hat{d}_{2\sigma}+\hat{d}_{3\sigma})+t^{\prime}\hat{d}^{\dagger}_{2\sigma}\hat{d}_{3\sigma}+\operatorname{{\rm H.c.}}\big{]}\\ &+V\sum_{\sigma}(\hat{d}_{1\sigma}^{\dagger}\hat{c}_{0\sigma}^{\phantom{\dagger}}+\hat{c}_{0\sigma}^{\dagger}\hat{d}_{1\sigma}^{\phantom{\dagger}})\;,\end{split}$ (53)
where $\hat{n}_{i}^{d}=\sum_{\sigma}d_{i\sigma}^{\dagger}d_{i\sigma}^{\phantom{\dagger}}$ is the number operator for each quantum dot, $\hat{N}^{d}=\sum_{i}\hat{n}^{d}_{i}$ is the total TQD number operator, and $\hat{S}_{z}^{d}=\tfrac{1}{2}\sum_{i}(d_{i\uparrow}^{\dagger}d_{i\uparrow}^{\phantom{\dagger}}-d_{i\downarrow}^{\dagger}d_{i\downarrow}^{\phantom{\dagger}})$ is the total TQD spin projection.
The UML analysis for the TQD proceeds similarly to that of the AIM. However, with the application of a magnetic field $B$, the spin multiplets previously labelled by the quantum number $S$ are now no longer $12S+$ 1-fold degenerate (although total $S_{z}$ is still conserved). Due to the broken particle-hole symmetry, $Q=+1$ and $-1$ sectors are no longer degenerate. The lower symmetry of the bare TQD model therefore implies that we also have spin and charge anisotropies in the MC effective model. As a consequence, UML for the TQD entails matching 8 independent observables, rather than 3 for the symmetric AIM. Thermal expectation values of the symmetry projectors $\hat{P}_{Q,S_{z}}$ are the observables which we compare in bare and effective models, with $(Q,S_{z})=(0,0)$, $(0,1)$, $(0,-1)$, $(1,1/2)$, $(1,-1/2)$, $(-1,1/2)$, $(-1,-1/2)$. The $(0,0)$ block is a diagonal $2\times 2$ matrix. These observables can be defined and computed (here using NRG) in both bare and effective models.
The effective model is then optimized by minimizing the KLD loss function by gradient descent. Since the KLD is still convex and its gradient, Eq. 10, is expressed in terms of the computed observables $\langle\hat{P}_{Q,S_{z}}\rangle$, the optimization remains highly efficient, despite the increase from 3 to 8 in the number of parameters to be tuned. The gradient descent method is easily able to treat optimization problems of this complexity [12]. Indeed, as shown later, systems with a far larger number of parameters in the non-convex case are also readily optimized.
Each step of the optimization process involves solving the MC effective model using the updated coupling constants, and computing the updated set of observables. The compute time simply grows linearly in the number of observables. These are compared with those of the bare model from a one-shot calculation. Even though the effective model has to be solved many times, these calculations are comparatively cheap, since by construction the Hilbert space dimension of the impurity in the effective model is less than that in the bare model. Although the effective model may have a larger number of coupling constants than the original bare model, it is defined in a reduced Hilbert space and therefore is of lower computational complexity to solve.
The iterative solution and refinement of the effective model with UML has essentially fixed computational cost, and is highly efficient. The potential bottleneck is in the one-shot solution of more complex bare models, as noted above. For the TQD however, NRG can still be used to obtain exact benchmark results.
We demonstrate the viability of UML for the TQD, Eq. LABEL:eq:tqd, in Fig. 6. Thermodynamic properties of the bare TQD (dot-dashed lines) and the optimized MC effective model (solid lines) are compared, for different temperatures $T$ and field strengths $B$. Panel (b) shows the impurity contribution to the total entropy, $S_{\rm imp}(T)$, and (c) shows the magnetic susceptibility $T\chi_{\rm imp}(T)$, both vs $T/B$. The simplified effective model successfully captures the physics of the TQD for $T\ll U_{d}$. At higher temperatures, TQD degrees of freedom not included in the effective model contribute to thermodynamic averages.
As $B$ is increased in the bare TQD model, we see a crossover from Kondo-screening of the TQD local moment, to polarization of the local moment by the field. Both the entropy and the magnetic susceptibility are fully quenched as $T\to 0$ for any $B$. However, the scale on which this happens depends on the underlying mechanism. For small fields $B\ll T_{K}$ (green line) screening occurs due to the Kondo effect, and the physics is insensitive to $B$. However, at larger $B$ the Kondo effect, which depends on dynamical spin-flip scattering, is destroyed. The TQD local moment is then screened on the scale of $B$. For $B\gg T_{K}$ we therefore see scaling collapse in terms of $T/B$, with polarization of the local moment setting in for $T/B\sim 1$ (see yellow, blue, red lines). The low-energy behavior of the optimized effective model captures this crossover and the onset of scaling vs $T/B$ accurately.
We conclude that more complex models, including those with more interacting degrees of freedom and lower symmetries, can be treated efficiently and accurately with UML. Physical properties of the resulting optimized effective models agree quantitatively at low temperatures with bare model benchmark reference results.
@@ 00 $\tfrac{3}{2}$ 0 ±${}\pm\tfrac{1}{2},~{}\pm\tfrac{3}{2}~{}$ $1$
If desired, the projectors $\hat{P}_{QST}$ can be formulated in second-quantized form (see Appendix D). New types of coupling term arise in the extended MC effective model, relative to the basic MC model. In particular, we see assisted hopping terms, in which an impurity spin-flip is accompanied by electronic tunneling between bath sites. This might be expected from the physical interpretation of the Kondo effect in terms of spin-flip scattering.
@@ IV.3 UML optimization of extended models
Figure 7: Analysis of UML optimization of the extended MC effective models (Eqs. 60 and 62), taking the AIM as our bare model. (a) Comparison of relative errors in computed observables (Eq. 64) as a function of the GD optimization parameter update step (epoch). Lines are color-coded according to the key, labelled by the (diagonal) projector quantum numbers $(Q,S,T)$. (b) Convergence of the Kondo temperature $T_{K}$ with epoch (red line) compared with the reference result in the bare model (blue line). (c) Comparison of relative (logarithmic) error in the Kondo temperature for the fixed-basis effective model Eq. 60 (pink line) and generalized basis model Eq. 62 (red line). Plotted for $0.38V^{2}/U_{d}=$ 0.3 and $U_{d}=0.4$, with the optimization performed at $T=0$ using NRG as the impurity solver for both bare and effective models.
@@
Taking the AIM again as our bare model for demonstration, we perform UML optimization on the extended MC effective model Eq. 62. For representative AIM parameters, in Fig. 7(a) we plot the relative error of selected (diagonal) observables labelled by the projector quantum numbers $(Q,S,T)$, as a function of the optimization step (epoch). We find that most effective model observables converge exponentially quickly towards their reference bare model counterparts, whereas a few are essentially stationary and cannot be matched. Indeed, it is in general not possible to perfectly match multiple observables simultaneously, and the minimum relative error is small but remains finite after convergence. This is because the effective model has good but not unlimited expressibility: minimization of the KLD loss function typically involves a compromise in the optimized solution. However, as Fig. 7(b) shows, this does not make much difference to the quality of the optimized effective model, as quantified by the Kondo temperature $T_{K}$. After $\sim 100$ or so training steps, $T_{K}$ for the effective model (red line) converges quite accurately to that of the bare model (blue line).